# A Survey on Text-Driven 360-Degree Panorama Generation

### v1

@@ Abstract
1 Introduction
@@
Rapid growth of immersive technologies, such as virtual reality (VR) and augmented reality (AR), has dramatically increased the demand for high-quality panoramic visual content. Among such content, 360-degree panoramas are pivotal in delivering realistic and immersive experiences by capturing a complete spherical view of an environment. Traditionally, producing these panoramas requires specialized camera equipment and considerable technical expertise. However, recent advances in text-driven 360-degree panorama generation Chen et al. (2022); Dastjerdi et al. (2022); Wang et al. (2024); Ye et al. (2024) have introduced groundbreaking capabilities, enabling the synthesis of 360-degree panoramic images directly from textual descriptions. This innovation can potentially revolutionize content creation over diverse domains including VR/AR applications, gaming, and virtual tours.
@@
Unlike conventional 2D images, 360-degree panoramas, often represented through equirectangular projection da Silveira et al. (2022), encompass the entire $360\degree\times 180\degree$ field of view (FoV). This distinctive format poses unique challenges for text-driven generation, requiring not only accurate image synthesis but also excellent preservation of geometric consistency and seamless visual coherence across the full $360\degree$ horizontal and $180\degree$ vertical extents.
@@
The availability of large-scale paired image-text datasets has facilitated the development of text-to-image latent diffusion models (LDMs) Rombach et al. (2022), which excel at synthesizing high-quality, visually compelling images aligned with given text descriptions. Leveraging the powerful generative capabilities of pre-trained LDMs, researchers have developed methods specifically tailored to address the unique challenges of text-driven 360-degree panorama generation Wang et al. (2023, 2024); Lu et al. (2024); Feng et al. (2023); Zhang et al. (2024). Despite the rapid progress in this field, there is currently no survey that provides a systematic analysis of these emerging methodologies, elucidates their key characteristics, and highlights their applications. To address this gap, this paper presents a holistic survey and analysis of the latest research advances in text-driven 360-degree panorama generation.
@@
This survey is structured as follows: First, we establish a foundational understanding of this field by introducing the principal representations of 360-degree panoramas, presenting prominent datasets commonly used in this area, and outlining key evaluation metrics employed to assess the quality and fidelity of generated panoramic content. Next, we review state-of-the-art methods for text-driven 360-degree panorama generation, categorizing them into two primary paradigms: (a) Text-Only Generation and (b) Text-Driven Narrow Field-of-View (NFoV) Outpainting. Methods on text-only generation aim to synthesize entire 360-degree panoramas exclusively from textual prompts, typically utilizing fine-tuning techniques such as Low-Rank Adaptation (LoRA) Hu et al. (2022) or DreamBooth Ruiz et al. (2023) on pre-trained LDMs Rombach et al. (2022). In contrast, text-driven NFoV outpainting allows greater user control by conditioning the generation process on both textual descriptions and initial NFoV images. Following this, we explore emerging applications of text-driven 360-degree panorama generation in the realm of 360-degree 3D scene generation. Finally, we discuss the prevailing challenges in this developing field and propose potential directions for future research.
@@
2 Related Work
2.1 Text-to-Image Diffusion Models
Text-to-image (T2I) diffusion models Nichol et al. (2022); Saharia et al. (2022); Ramesh et al. (2022); Rombach et al. (2022); Podell et al. (2024) have achieved remarkable progress in generating high-fidelity and photorealistic images from textual descriptions. These models have garnered widespread attention because of their intuitive text-based conditioning as a user-friendly interface for diverse image generation tasks.
T2I diffusion models can be broadly categorized into pixel-space and latent-space models. Pixel-space models, such as GLIDE Nichol et al. (2022) and Imagen Saharia et al. (2022), operate directly in the pixel space, producing visually impressive results at the expense of substantial computational resources, limiting their scalability. In contrast, latent diffusion models (LDMs) Rombach et al. (2022) address these limitations by leveraging pre-trained autoencoders like VQGAN Esser et al. (2021) to map images into a compact latent space, where the diffusion process is conducted. This reduces computational overhead while maintaining high-quality outputs, making LDMs a preferred framework for text-driven 360-degree panorama generation, as surveyed in this work.
2.2 3D Scene Representation
@@ 2.1 Text-to-Image Diffusion Models
Neural implicit functions Park et al. (2019); Mescheder et al. (2019); Xia and Xue (2023), which represent 3D scenes as continuous functions encoded within neural network parameters, offer a compact and flexible paradigm for scene representation. Notably, Neural Radiance Fields (NeRFs) Mildenhall et al. (2020) stand out for their ability to achieve high-quality novel view synthesis. However, NeRF’s reliance on dense volumetric sampling along camera rays results in slow training, hindering its practicability.
@@
Recently, 3D Gaussian Splatting (3DGS) Kerbl et al. (2023) has emerged as an efficient alternative to 3D scene representation. By combining an explicit representation of 3D Gaussians with a highly efficient differentiable rasterization pipeline, 3DGS facilitates rapid scene reconstruction and rendering. This advancement has opened up new possibilities, including recent explorations in text-to-3D 360-degree scene synthesis Zhou et al. (2024b); Li et al. (2024), which leverage text-driven 360-degree panorama generation techniques.
3 Preliminaries
3.1 Representations of 360-Degree Panoramas
@@
The representation of 360-degree panoramic content poses a fundamental challenge: How to accurately map spherical visual information onto a two-dimensional plane? To address this, a variety of projection methodologies da Silveira et al. (2022) have been developed, each with distinct advantages and trade-offs. Below, we outline two widely used formats for 360-degree panorama generation: Equirectangular Projection and Cubemap Projection.
Equirectangular Projection (ERP)
@@
As the most prevalent representation format for 360-degree panoramas, ERP establishes a direct mapping between spherical and planar coordinates: longitude corresponds to the horizontal axis, spanning the full 360°$\degree$ range, while latitude maps to the vertical axis, covering 180°$\degree$ from -90°$\degree$ (south pole) to +90°$\degree$ (north pole). ERP’s simplicity and compatibility with web viewers and VR headsets make it the preferred choice for numerous applications. Additionally, its representation as a single, continuous image allows the direct application of image manipulation techniques, such as text-driven 360-degree panorama-to-panorama translation Wang and Xue (2025). Despite these advantages, ERP introduces pronounced geometric distortions, particularly at the polar regions, where the visual content appears stretched. Furthermore, the texel density of a spherical image in ERP is non-uniformly distributed: it is comparatively lower in the equatorial regions and markedly higher towards the poles. This inhomogeneity can be particularly problematic in scenarios where critical visual information is predominantly located away from the poles, leading to inefficient utilization of image resolution for regions of interest. In this survey, unless explicitly stated otherwise, 360-degree panoramas are represented using ERP.
Cubemap Projection (CMP)
@@
CMP offers an alternative representation that mitigates the distortions inherent in the ERP format, particularly at the poles. In CMP, the spherical image is projected onto the six faces of a cube, with each face representing a 90°$\degree$$\times$90°$\degree$ field of view. This division significantly reduces geometric distortions, making CMP more compatible with diffusion priors from text-to-image diffusion models trained on standard perspective images Kalischek et al. (2025). However, CMP introduces several challenges: (1) it increases the complexity of image manipulation compared to the single-image format of ERP; (2) it may necessitate additional conversion for compatibility with platforms or viewers that primarily support ERP. Despite these practical challenges, CMP is well-suited for applications that demand reduced distortion and higher fidelity. The width and height of an ERP image are four and two times the side length of the corresponding CMP, respectively, reflecting the geometric relationship between the two formats.
3.2 Datasets
360-degree panoramic image generation from text prompts presents unique challenges due to the complete 360°$\degree$$\times$180°$\degree$ field of view that these images encompass. Text-to-image diffusion models Saharia et al. (2022); Rombach et al. (2022); Podell et al. (2024), predominantly trained on perspective images with a narrower field of view, often struggle to synthesize high-quality 360-degree panoramas. To address this, several specialized datasets have been developed to facilitate research in this domain. Tab. 1 summarizes these datasets, with further details provided below.
@@ 3.2 Datasets
Table 1: Summary of Text-Driven Panorama Generation Datasets: Includes categories, publication year, sample size, resolution (Res.), and license. Categories are indoor (I), outdoor (O), or hybrid (I, O). Datasets marked with $\star$ are sourced from public websites.
SUN360 Xiao et al. (2012) is a comprehensive database comprising 67,583 high-resolution (9104$\times$4552) panoramic images sourced from the Internet. Each image covers a 360°$\degree$$\times$180°$\degree$ field of view in ERP format and is manually categorized into 80 distinct classes. Originally created for scene viewpoint recognition, SUN360 now serves as a valuable resource for a wide range of computer vision, computer graphics, and related research areas.
Matterport3D Chang et al. (2017) offers 10,800 indoor 360-degree panoramic images with corresponding depth maps, all at a resolution of 2048$\times$1024 pixels. These panoramas are derived from 194,400 RGB-D images of 90 buildings, making it a rich dataset for studying indoor environments.
Laval Indoor Gardner et al. (2017) consists of 2,233 high-dynamic-range, high-resolution (7768$\times$3884) 360-degree panoramic images, specifically curated for the study of extensive indoor scenes, such as factories, apartments, and houses.
Laval Outdoor Hold-Geoffroy et al. (2019) complements its indoor counterpart, offering 205 high-dynamic-range, high-resolution (7768$\times$3884) 360-degree panoramic images that capture diverse outdoor environments, including urban and natural scenes.
Structured3D Zheng et al. (2020) contains 196,515 360-degree panoramas with varying configurations and lighting conditions, representing 21,835 distinct rooms. Rendered at a resolution of 1024$\times$512 from 3D scenes of original house design files, Structured3D is ideal for research on structured 3D modeling and understanding.
Pano360 Kocabas et al. (2021) contains 35,000 360-degree panoramic images with a resolution of 8192$\times$4096. Of these, 34,000 are sourced from Flickr, with the remainder rendered from photorealistic 3D scenes. Pano360 was originally proposed for training camera calibration networks.
Polyhaven Poly Haven (2025) contributes 786 real-world high-resolution (8192$\times$4096) 360-degree panoramas encompassing a variety of indoor and outdoor scenes.
Humus Persson (2025) includes 139 real-world 360-degree panoramas represented using cubemap projection, with each face having a resolution of 2048$\times$2048 pixels. This dataset includes indoor and outdoor environments.
Figure 1: Paradigms for Text-Driven 360-Degree Panorama Generation. (a) Text-Only Generation synthesizes 360-degree panoramas from textual descriptions only. (b) Text-Driven NFoV Outpainting uses prompts and initial NFoV images as input to generate panoramic images.
3.3 Evaluation Metrics
@@
FID Heusel et al. (2017) measures the distance between feature distributions of generated and real images using a pre-trained Inception-v3 network Szegedy et al. (2016). Lower FID scores indicate better perceptual quality and closer alignment with the real image distribution.
@@
KID Bińkowski et al. (2018) measures the difference between real and generated image distributions by computing the maximum mean discrepancy of their features extracted from Inception-v3 Szegedy et al. (2016). Similar to FID, lower KID values indicate better image quality.
@@
IS Salimans et al. (2016) measures both the quality and diversity of generated images by leveraging Inception-v3 Szegedy et al. (2016). It calculates the KL divergence between the conditional class distribution of generated images and the marginal distribution over all generated samples. Higher IS scores suggest better visual quality and diversity.
@@
CS Radford et al. (2021) evaluates consistency between text prompts and generated images using the CLIP model Radford et al. (2021). It calculates the cosine similarity between the text embedding of the prompt and the visual embedding of the generated image. A higher CS score reflects stronger text-image alignment and semantic coherence.
The universal metrics provide general assessments but often fail to capture nuances and geometric distortions unique to 360-degree panoramas. Panorama-specific metrics designed specifically for ERP-based 360-degree panoramic images include Fréchet Auto-Encoder Distance (FAED), Omnidirectional FID (OmniFID), and Discontinuity Score (DS).
FAED Oh et al. (2022) computes Fréchet distances between features extracted from generated and real panoramas. Unlike FID, it employs an autoencoder Hinton and Salakhutdinov (2006) specifically trained on 360-degree panoramic images. Lower FAED scores reflect better perceptual and geometric quality tailored to the unique panoramic properties.
OmniFID Christensen et al. (2024) adapts FID to specifically evaluate 360-degree panoramas. It converts equirectangular panoramas into cubemap representations and calculates FID across three disjoint subsets of cubemap faces, averaging the results. Lower OmniFID scores indicate higher geometric fidelity in 360-degree panorama generation.
DS Christensen et al. (2024) measures the seam alignment across the borders of generated panoramas by applying a kernel-based edge detection algorithm. Lower discontinuity scores correspond to fewer visible seam artifacts, indicating better perceived consistencies across the seam.
We provide a comprehensive comparison of state-of-the-art methods, introduced in the following section, using the outlined metrics in Sec. 4.3.
4 State-of-the-Art Methods
Existing text-driven 360-degree panorama generation methods can be broadly categorized into two paradigms according to input modalities: (a) Text-Only Generation aims to synthesize 360-degree panoramas from textual prompts only, while (b) Text-Driven NFoV Outpainting leverages both textual descriptions and initial NFoV images to guide the generation process, offering enhanced user control. Fig. 1 provides an intuitive illustration for both paradigms. We detail the literature for both as follows.
4.1 Text-Only Generation
This paradigm focuses on synthesizing 360-degree panoramas from textual descriptions. Tab. 2 provides a comparative overview of representative text-only methods.
Text2Light Chen et al. (2022), an early notable effort, explored a hierarchical framework utilizing VQGAN Esser et al. (2021) and CLIP Radford et al. (2021) to address this challenge. Recently, the advent of latent diffusion models (LDMs) Rombach et al. (2022) for text-to-image synthesis marked a significant advancement, enabling more sophisticated 360-degree panorama generation techniques. LDMs are typically trained on vast datasets consisting of perspective images with a limited field of view and corresponding text descriptions. Despite demonstrating robust capabilities in generating perspective images from text prompts, these models face significant difficulties when creating 360-degree panoramas with a complete 360°$\degree$$\times$180°$\degree$ field of view, which differ substantially from traditional perspective images.
Fine-Tuning.
To adapt pre-trained LDMs for 360-degree panorama synthesis, a common strategy is to fine-tune these models with specialized 360-degree panorama datasets. Diffusion360 Feng et al. (2023) exemplifies this approach by leveraging the DreamBooth technique Ruiz et al. (2023) to fine-tune a pre-trained LDM Rombach et al. (2022) on SUN360 Xiao et al. (2012). To ensure geometric consistency of boundaries, Diffusion360 utilizes a circular blending strategy during both the denoising process and the VAE decoding stage, effectively reducing seam artifacts. In addition, it introduces a super-resolution module to enable the generation of high-resolution (6114$\times$3072) 360-degree panoramas. Notably, the DreamBooth fine-tuning approach is computationally demanding and time-consuming because it generally fine-tunes a significant portion of the model parameters.
In contrast, LoRA Hu et al. (2022) has recently gained attention as a parameter-efficient fine-tuning method. LoRA works by injecting trainable low-rank matrices into the pre-trained model’s weights, allowing for rapid adaptation to new tasks with minimal additional parameters. For example, StitchDiffusion Wang et al. (2024) employs LoRA to fine-tune a pre-trained LDM on a dataset of paired 360-degree images (sourced from Polyhaven Poly Haven (2025)) and corresponding textual descriptions generated using BLIP Li et al. (2022). It further introduces a tailored generation method based on MultiDiffusion Bar-Tal et al. (2023) to ensure boundary continuity in generated panoramas.
Other works have similarly adopted the LoRA fine-tuning technique. PanFusion Zhang et al. (2024) devises a dual-branch diffusion model, trained on Matterport3D Chang et al. (2017), with separate LoRA layers to integrate both global panoramic and local perspective views. It introduces an equirectangular-perspective projection attention module to facilitate information exchange between the two branches, aiming to alleviate visual inconsistencies in the generated 360-degree panoramas. However, PanFusion’s output often exhibits blurriness at the top and bottom regions, due to its training dataset. To avoid this issue, DiffPano Ye et al. (2024) utilizes the Habitat Matterport 3D dataset Ramakrishnan et al. (2021) to produce multi-view consistent 360-degree panoramas with clearer top and bottom details. For generating more precise textual descriptions, DiffPano adopts BLIP2 Li et al. (2023) and Llama2 Touvron et al. (2023) sequentially, resulting in a panoramic video-text dataset. Based on this dataset, it fine-tunes a pre-trained LDM using LoRA for single-view text-driven 360-degree panorama generation. Furthermore, to enable multi-view 360-degree panorama generation based on text prompts and camera viewpoints, DiffPano proposes a spherical epipolar-aware attention module, which effectively extends single-view panorama generation to multi-view consistent panorama generation.
Training-Free.
PanoFree Liu et al. (2024), on the other hand, departs from fine-tuning approaches by introducing a tuning-free multi-view image generation framework based on a pre-trained LDM. Guided by textual descriptions, PanoFree leverages iterative warping and inpainting steps to produce multi-view perspective images, which are subsequently stitched into 360-degree panoramas, thus avoiding the need for specialized 360-degree panorama datasets.
Method Publication LDM-B TF Code
Text2Light TOG 2022 $\times$ $\times$ $\checkmark$
Diffusion360 arxiv 2023 $\checkmark$ $\times$ $\checkmark$
StitchDiffusion WACV 2024 $\checkmark$ $\times$ $\checkmark$
PanFusion CVPR 2024 $\checkmark$ $\times$ $\checkmark$
PanoFree ECCV 2024 $\checkmark$ $\checkmark$ $\times$
DiffPano NeurIPS 2024 $\checkmark$ $\times$ $\times$

### v2

@@ Abstract
Index Terms:
360-degree panorama generation, text-driven generation, text-to-image diffusion model, 360-degree 3D scene generation.
I Introduction
@@
Rapid growth of immersive technologies, such as virtual reality (VR) and augmented reality (AR), has dramatically increased the demand for high-quality panoramic visual content. Among such content, 360-degree panoramas are pivotal in delivering realistic and immersive experiences by capturing a complete spherical view of an environment. Traditionally, producing these panoramas requires specialized camera equipment and considerable technical expertise. However, recent advances in text-driven 360-degree panorama generation [14, 40, 56, 57] have introduced groundbreaking capabilities, enabling the synthesis of 360-degree panoramic images directly from textual descriptions. This innovation can potentially revolutionize content creation over diverse domains including VR/AR applications, gaming, and virtual tours.
@@
Unlike conventional 2D images, 360-degree panoramic images, often represented through equirectangular projection [7], encompass the entire 360°$\degree$$\times$180°$\degree$ field of view, as shown in Fig. 1. This distinctive format poses unique challenges for text-driven generation, requiring not only accurate image synthesis but also excellent preservation of geometric consistency and seamless visual coherence across the full $360\degree$ horizontal and $180\degree$ vertical extents.
Figure 1: Visual comparison between a 360-degree panoramic image and a conventional 2D image.
@@
The availability of large-scale paired image-text datasets has facilitated the development of text-to-image latent diffusion models (LDMs) [63], which excel at synthesizing high-quality, visually compelling images aligned with given text descriptions [64, 65, 66, 67, 72]. Leveraging the powerful generative capabilities of pre-trained LDMs, researchers have developed methods specifically tailored to address the unique challenges of text-driven 360-degree panorama generation [42, 57, 58, 59, 60, 73]. Despite the rapid progress in this field, there is currently no survey that provides a systematic analysis of these emerging methodologies, elucidates their key characteristics, and highlights their applications. To address this gap, this paper presents a holistic survey and analysis of the latest research advances in text-driven 360-degree panorama generation.
@@
This survey is structured as follows: First, we establish a foundational understanding of this field by introducing the principal representations of 360-degree panoramas, presenting prominent datasets commonly used in this area, and outlining key evaluation metrics employed to assess the quality and fidelity of generated panoramic content. Next, we review state-of-the-art (SOTA) methods for text-driven 360-degree panorama generation, categorizing them into two primary paradigms: (a) Text-Only Generation and (b) Text-Driven Narrow Field-of-View (NFoV) Outpainting. Fig. 2 and Fig. 3 provide a systematic taxonomy and a chronological overview of these SOTA methods, respectively. Methods on text-only generation aim to synthesize entire 360-degree panoramas exclusively from textual prompts, typically utilizing fine-tuning techniques such as Low-Rank Adaptation (LoRA) [39] or DreamBooth [38] on pre-trained LDMs [63]. In contrast, text-driven NFoV outpainting allows greater user control by conditioning the generation process on both textual descriptions and initial NFoV images. Following this, we explore emerging applications of text-driven 360-degree panorama generation in the realm of 360-degree 3D scene generation. Finally, we discuss the prevailing challenges in this developing field and propose potential directions for future research.
Figure 2: A systematic taxonomy proposed in this survey of text-driven 360-degree panorama generation methods.
Figure 3: Chronological overview of text-driven 360-degree panorama generation approaches. Methods in lime, orange, violet, and cyan are from Sec. IV-A1, Sec. IV-A2, Sec. IV-B1, and Sec. IV-B2, respectively.
@@
II Related Work
II-A Text-to-Image Diffusion Models
Text-to-image (T2I) diffusion models [11, 12, 13, 21, 63] have achieved remarkable progress in generating high-fidelity and photorealistic images from textual descriptions. These models have garnered widespread attention because of their intuitive text-based conditioning as a user-friendly interface for diverse image generation tasks.
T2I diffusion models can be broadly categorized into pixel-space and latent-space models. Pixel-space models, such as GLIDE [11] and Imagen [12], operate directly in the pixel space, producing visually impressive results at the expense of substantial computational resources, limiting their scalability. In contrast, latent diffusion models (LDMs) [63] address these limitations by leveraging pre-trained autoencoders like VQGAN [43] to map images into a compact latent space, where the diffusion process is conducted. This reduces computational overhead while maintaining high-quality outputs, making LDMs a preferred framework for text-driven 360-degree panorama generation, as surveyed in this work.
II-B 3D Scene Representation
@@ II-A Text-to-Image Diffusion Models
Neural implicit functions [9, 10, 29], which represent 3D scenes as continuous functions encoded within neural network parameters, offer a compact and flexible paradigm for scene representation. Notably, Neural Radiance Fields (NeRFs) [8] stand out for their ability to achieve high-quality novel view synthesis. However, NeRF’s reliance on dense volumetric sampling along camera rays results in slow training, hindering its practicability.
@@
Recently, 3D Gaussian Splatting (3DGS) [26, 35] has emerged as an efficient alternative to 3D scene representation. By combining an explicit representation of 3D Gaussians with a highly efficient differentiable rasterization pipeline, 3DGS facilitates rapid scene reconstruction and rendering. This advancement has opened up new possibilities, including recent explorations in text-to-3D 360-degree scene synthesis [33, 32], which leverage text-driven 360-degree panorama generation techniques.
III Preliminaries
III-A Representations of 360-Degree Panoramas
@@
The representation of 360-degree panoramic content poses a fundamental challenge: How to accurately map spherical visual information onto a two-dimensional plane? To address this, a variety of projection methodologies [7] have been developed, each with distinct advantages and trade-offs. Below, we first outline two widely used formats for 360-degree panorama representation: Equirectangular Projection (ERP) and Cubemap Projection (CMP), as illustrated in Fig. 4, and then introduce Multi-Perspective Projection (MPP), a category we define for the classification purposes of this survey.
Figure 4: Visual comparison between the equirectangular and cubemap projections of spherical images (360-degree panoramic images).
III-A1 Equirectangular Projection (ERP)
@@
As the most prevalent representation format for 360-degree panoramas, ERP establishes a direct mapping between spherical and planar coordinates: longitude corresponds to the horizontal axis, spanning the full 360°$\degree$ range, while latitude maps to the vertical axis, covering 180°$\degree$ from -90°$\degree$ (south pole) to +90°$\degree$ (north pole). ERP’s simplicity and compatibility with web viewers and VR headsets make it the preferred choice for numerous applications. Additionally, its representation as a single, continuous image allows the direct application of image manipulation techniques, such as text-driven 360-degree panorama-to-panorama translation [28]. Despite these advantages, ERP introduces pronounced geometric distortions, particularly at the polar regions, where the visual content appears stretched. Furthermore, the texel density of a spherical image in ERP is non-uniformly distributed: it is comparatively lower in the equatorial regions and markedly higher towards the poles. This inhomogeneity can be particularly problematic in scenarios where critical visual information is predominantly located away from the poles, leading to inefficient utilization of image resolution for regions of interest. In this survey, unless explicitly stated otherwise, 360-degree panoramas are represented using ERP.
III-A2 Cubemap Projection (CMP)
@@
CMP offers an alternative representation that mitigates the distortions inherent in the ERP format, particularly at the poles. In CMP, the spherical image is projected onto the six faces of a cube, with each face representing a 90°$\degree$$\times$90°$\degree$ field of view. This division significantly reduces geometric distortions, making CMP more compatible with diffusion priors from text-to-image diffusion models trained on standard perspective images [44]. However, CMP introduces several challenges: (1) it increases the complexity of image manipulation compared to the single-image format of ERP; (2) it may necessitate additional conversion for compatibility with platforms or viewers that primarily support ERP. Despite these practical challenges, CMP is well-suited for applications that demand reduced distortion and higher fidelity. The width and height of an ERP image are four and two times the side length of the corresponding CMP, respectively, reflecting the geometric relationship between the two formats.
III-A3 Multi-Perspective Projection (MPP)
MPP is defined as the use of multiple individual perspective images to collectively represent a 360-degree panoramic view, where these images may or may not overlap. This category is characterized by configurations that deviate from the standard six-face, 90°$\degree$$\times$90°$\degree$ field of view CMP.
TABLE I: Summary of Popular Datasets for Text-Driven 360-Degree Panorama Generation: Includes categories, publication year, sample size, resolution (Res.), and license. Categories are indoor (I), outdoor (O), or hybrid (I, O). Datasets marked with $\star$ are sourced from public websites.
@@ III-A3 Multi-Perspective Projection (MPP)
III-B Datasets
360-degree panoramic image generation from text prompts presents unique challenges due to the complete 360°$\degree$$\times$180°$\degree$ field of view that these images encompass. Text-to-image diffusion models [12, 21, 63], predominantly trained on perspective images with a narrower field of view, often struggle to synthesize high-quality 360-degree panoramas. To address this, several specialized datasets have been developed to facilitate research in this domain. Tab. I summarizes these datasets, with further details provided below.
SUN360 [48] is a comprehensive database comprising 67,583 high-resolution (9104$\times$4552) panoramic images sourced from the Internet. Each image covers a 360°$\degree$$\times$180°$\degree$ field of view in ERP format and is manually categorized into 80 distinct classes. Originally created for scene viewpoint recognition, SUN360 now serves as a valuable resource for a wide range of computer vision, computer graphics, and related research areas.
Matterport3D [49] offers 10,800 indoor 360-degree panoramic images with corresponding depth maps, all at a resolution of 2048$\times$1024 pixels. These panoramas are derived from 194,400 RGB-D images of 90 buildings, making it a rich dataset for studying indoor environments.
Laval Indoor [45] consists of 2,233 high-dynamic-range, high-resolution (7768$\times$3884) 360-degree panoramic images, specifically curated for the study of extensive indoor scenes, such as factories, apartments, and houses.
Laval Outdoor [46] complements its indoor counterpart, offering 205 high-dynamic-range, high-resolution (7768$\times$3884) 360-degree panoramic images that capture diverse outdoor environments, including urban and natural scenes.
Structured3D [47] contains 196,515 360-degree panoramas with varying configurations and lighting conditions, representing 21,835 distinct rooms. Rendered at a resolution of 1024$\times$512 from 3D scenes of original house design files, Structured3D is ideal for research on structured 3D modeling and understanding.
Pano360 [19] contains 35,000 360-degree panoramic images with a resolution of 8192$\times$4096. Of these, 34,000 are sourced from Flickr, with the remainder rendered from photorealistic 3D scenes. Pano360 was originally proposed for training camera calibration networks.
Polyhaven [5] contributes 786 real-world high-resolution (8192$\times$4096) 360-degree panoramas encompassing a variety of indoor and outdoor scenes.
Humus [6] includes 139 real-world 360-degree panoramas represented using cubemap projection, with each face having a resolution of 2048$\times$2048 pixels. This dataset includes indoor and outdoor environments.
Figure 5: Paradigms for Text-Driven 360-Degree Panorama Generation. (a) Text-Only Generation synthesizes 360-degree panoramas from textual descriptions only. (b) Text-Driven NFoV Outpainting uses prompts and initial NFoV images as input to generate 360-degree panoramic images.
III-C Evaluation Metrics
@@
FID [53] measures the distance between feature distributions of generated and real images using a pre-trained Inception-v3 network [36]. Lower FID scores indicate better perceptual quality and closer alignment with the real image distribution.
@@
KID [20] measures the difference between real and generated image distributions by computing the maximum mean discrepancy of their features extracted from Inception-v3 [36]. Similar to FID, lower KID values indicate better image quality.
@@
IS [54] measures both the quality and diversity of generated images by leveraging Inception-v3 [36]. It calculates the KL divergence between the conditional class distribution of generated images and the marginal distribution over all generated samples. Higher IS scores suggest better visual quality and diversity.
@@
CS [52] evaluates consistency between text prompts and generated images using the CLIP model [52]. It calculates the cosine similarity between the text embedding of the prompt and the visual embedding of the generated image. A higher CS score reflects stronger text-image alignment and semantic coherence.